# Choi-Defined Resource Theories

### v1

@@ Abstract
The resource theories of separable entanglement, non-positive partial transpose entanglement, magic, and imaginarity share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state. In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories. We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and we prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations.
@@
The recent booming interest in quantum computing, communication, and information originates from the understanding that quantumness allows us to outperform the classical world, which means that quantum objects are resources. The new technological developments, which are now becoming commercially available, demand a rigorous and thorough understanding of quantum resources, in order to quantify them precisely, and optimize their use. This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Resource theories are based on the simple observation that agents in a laboratory can only perform a subset of all possible deterministic quantum operations due to the various constraints on the agents and the systems involved. For instance, when two agents, Alice and Bob, are physically separated, they may not have a reliable means to exchange quantum systems, limiting them to local operations and classical communication (LOCC) [5, 6, 7]. Such allowed operations are called free. While free states are typically defined as those states that can be prepared with free operations, there are situations where it is beneficial to construct free operations starting from a set of free states. As an example, Alice and Bob may quickly realize that they can create all and only separable states and wonder what free operations are compatible with their set of free states. There are different ways to construct free operations starting from free states [1]. However, the minimal condition is that one cannot generate a resource by applying a free operation on a free state in a complete sense, i.e., even if applied only to part of a bipartite state. Operations defined with such a construction are known as completely resource non-generating (CRNG) operations [1]. CRNG operations are of particular interest because they are the largest set of free operations compatible with a given set of free states, and they can be used to derive lower or upper bounds in any resource theory with the same set of free states. For example, separable and positive partial transpose operations have been used to overcome the challenging characterization of LOCC operations [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. These extensions of the resource theory of entanglement are known as the resource theories of separable entanglement (SEP) and non positive partial transpose entanglement (NPT), respectively. In other settings, such as in the resource theory of magic [28, 29, 30, 31, 32] and imaginarity [33, 34, 35] the CRNG operations are even more relevant because they coincide with the free operations [31, 32, 35]. Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state [37, 38, 46, 32, 35]. We call resource theories that exhibit this property Choi-defined resource theories. This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization [37, 38, 32, 35]. Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states. In this letter, we answer this question by providing necessary and sufficient conditions on the set of free states. Moreover, we show that when such a construction is possible, the resulting set of free operations coincides with the set of CRNG operations associated with those free states. These results generalize what has been already observed in the resource theories of magic, imaginarity, NPT, and SEP [37, 38, 46, 31, 32, 35].
@@
One of the constraints of the resource theory of bipartite entanglement is that the two physically separated agents cannot exchange quantum systems. This may appear to conflict with the condition that the swap channel is a free operation. However, in this scenario, the systems in question are bipartite systems [2], i.e., $A=A_{1}A_{2}$ and $B=B_{1}B_{2}$. Here the physical separation is between subscript $1$ and $2$. The swap channel simply swaps the order of the two bipartite systems, for example, $A_{1}A_{2}B_{1}B_{2}\leftrightarrow B_{1}B_{2}A_{1}A_{2}$. If we organize the systems according to the physical separation, the action of the swap channel is $A_{1}B_{1}A_{2}B_{2}\to B_{1}A_{1}B_{2}A_{2}$. Therefore, the two agents are locally exchanging the order of their own quantum systems and there is no exchange of quantum systems between them.
@@
Note that here one needs to know which system is the input of the original channel and which is the output. Clearly, if such information is missing, a matrix $M_{BA}$ could be associated with different linear maps, see the Supplemental Material (SM) for examples. Observe that, if we divide the Choi matrix $M_{B\mathord{:}A^{\prime}}$ of the channel $\mathcal{M}_{A\to B}$ by the dimension $d_{A}$ of $A$, we obtain a quantum state $\mu_{B\mathord{:}A^{\prime}}$ such that
@@
It is a natural question to characterize when, given a set of free states, the construction of a CDRT is allowed. In every well-defined resource theory, the identity is a free operations. As a consequence, its renormalized Choi matrix $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is free in every CDRT. Requesting that resource theories are closed under the sequential composition of free channels (Figure 1), we find a less trivial condition for the set free states. In Ref. [47, 48], the authors introduced an operation between bipartite matrices, called link product, defined as
@@
where $\cdot^{T_{B}}$ is the partial transpose on system $B$. This operation translates sequential composition of channels into an operation between Choi matrices. Thanks to the link product, we can restate the golden rule of resource theories as a condition on the set of free states. That is, the condition that $\mathcal{M}_{A\to B}(\rho_{A})$ be free whenever $\mathcal{M}_{A\to B}$ and $\rho_{A}$ are free becomes $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a free state if $\rho_{A}$ and $\mu_{B\mathord{:}A^{\prime}}$ are free states and $\mu_{B\mathord{:}A^{\prime}}$ is the renormalized Choi matrix of a quantum channel. We now state our first main result.
@@
Notably, this theorem provides an easy construction for CRNG operations when the set of free states satisfies the conditions of Theorem 3. For example, the set of separable states satisfies the conditions of Theorem 3, as we show later. Therefore, Theorems 3 and 5 imply that the CRNG operations are those that have a separable Choi matrix, i.e., the separable operations. Note that in general CRNG operations are not CD operations. Indeed, the CRNG operations in the resource theory of athermality are the Gibbs-preserving operations (see, e.g., Ref. [51]), but their renormalized Choi matrices are not free states.
§ V.  Example: resource theory of magic
Now we give a closer look at the resource theory of magic [28, 29, 30, 31, 32]. It originates from the observation that it is possible to achieve universal quantum computation by using only Clifford unitaries and some special states called magic states [52]. Being grounded in quantum computation, systems are usually systems of $n$ qubits [30]. In this scenario, Clifford unitaries, state preparation and measurements in the computational basis are easy to perform, and therefore free. In particular, Clifford unitaries are those transformations that can be implemented with circuits containing only Hadamard, CNOT, and phase-shift gates [53]. The free operations are called stabilizer operations, and analogously free states are called stabilizer states. The resources are magic states and operations.
We prove now that the set of magic states satisfies the conditions of Theorem 3. For the first condition, we denote the number of qubits in $A$ with $n$, and the elements of the computational basis with |$\Ket{x_{1}\dots x_{n}}$, where $x_{j}=0,1$. The normalized Choi state |$\Ket{\phi}=\frac{1}{2^{n/2}}\sum_{x_{1}\dots x_{n}}\Ket{x_{1}\dots x_{n}}\Ket{x_{1^{\prime}}\dots x_{n^{\prime}}}$ is obtained by preparing the $n$ entangled states $\frac{1}{\sqrt{2}}\sum_{x_{j}}\Ket{x_{j}x_{j^{\prime}}}$ and by swapping the order of the systems. The $j$-th 2-qubit entangled state can be obtained with an Hadamard and a CNOT gate acting on |$\Ket{0_{j}0_{j^{\prime}}}$, and the swap channels can be decomposed into three CNOT gates (see, e.g., Ref. [54]). All these operations are stabilizer operations, therefore the renormalized Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is a stabilizer state. This proves that condition 1 of Theorem 3 is satisfied. We now consider the closure under link product of the set of stabilizer states. Let $\mu_{BA^{\prime}}$ and $\rho_{A}$ be free states, where $A$ is a $n$-qubit system and $B$ is a $m$-qubit system. This implies that there exists a protocol to create $\mu_{BA^{\prime}}\otimes{\rho_{A}}$ for free. By Eq. (3), the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is obtained by a protocol with postselection:
  1. 1.
    Prepare $\mu_{BA^{\prime}}\otimes\rho_{A}$.
  2. 2.
    Perform a measurement in the Bell basis for each of the pair of qubits in $A^{\prime}A$
  3. 3.
    Postselect on the measurement outcome associated with $\frac{1}{d_{A}}\Phi_{A^{\prime}A}$. After the postselection, the system $B$ is in the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$, as requested.
Each of the steps of this protocol is a free operation. Indeed, the measurement in the Bell basis can be converted into a measurement in the computational basis by adding Hadamard and CNOT gates and these are stabilizer operations. Therefore the state $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is free, as requested by Theorem 3. The set of stabilizer states satisfy all the conditions of Theorem 3, and the CDRT constructed form it is the resource theory of magic [32]. Note that, by Theorem 5, in the resource theory of magic the free operations coincide with the CRNG operations. Note that Corollary 4 allows us to predict that the maximally mixed $n$-qubit state $\frac{1}{2^{n}}\operatorname{\mathds{1}}_{2}\otimes\dots\otimes\operatorname{\mathds{1}}_{2}$ is free. We can see this directly from the definition of stabilizer states. Indeed, each of the qubits is prepared as a probabilistic mixture of elements of the computational basis, and any state preparation in the computational basis is a free operation.
§ VI.  Example: resource theory of imaginarity
We now turn our attention to the resource theory of imaginarity [33, 34, 35]. The aim of this resource theory is to evaluate the role of imaginarity in quantum mechanics. A state is free if its density matrix, expressed with respect to a fixed basis, is real. States with density matrices containing non-real numbers are resources. The Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ relative to the fixed basis, has only real coefficients, and therefore is free. This proves condition 1 of Theorem 3. Condition 2 follows trivially because it can be expressed using only multiplications and traces of real matrices, from which it is possible to obtain only real matrices. This proves that it is possible to construct a CDRT from the set of real matrices, and it coincides with the resource theory of imaginarity, where the free operations are all the CRNG operations [35]. Note that, the maximally mixed state is real too, and therefore free, as predicted by Corollary 4.
§ VII.  Example: resource theory of separable entanglement
In the theory of entanglement [5, 6, 7, 8, 9, 10] systems are bipartite systems, and free states are separable states. In other words, each of the systems $A$ and $B$ mentioned in Theorem 3 is bipartite, i.e., $A=A_{1}A_{2}$, $B=B_{1}B_{2}$, and the physical separation is between systems with ‘$1$’ as subscript and systems with ‘$2$’ as subscript. To prove the first condition of Theorem 3, we fix an ortonormal basis for each system, e.g., {|$\Set{\Ket{a_{1}}}$ is a basis for $A_{1}$, and we write the Choi state as
$\Phi_{AA^{\prime}}$ $=\sum_{a_{1},a_{2}}|a_{1}a_{2}\rangle\!\langle a^{\prime}_{1}a^{\prime}_{2}|$ (5)
$=\sum_{a_{1}}|a_{1}\rangle\!\langle a^{\prime}_{1}|\otimes\sum_{a_{2}}|a_{2}\rangle\!\langle a^{\prime}_{2}|.$
Therefore $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is separable, free, and condition 1 of Theorem 3 is satisfied. We now prove that the set of separable state is closed under link product. Let $\mu_{BA^{\prime}}$ be a separable state, i.e., there exists density matrices $\mu^{j}_{B_{1}A^{\prime}_{1}}$ and $\tilde{\mu}^{j}_{B_{2}A^{\prime}_{2}}$ such that $\mu_{B_{1}A^{\prime}_{1}B_{2}A^{\prime}_{2}}=\sum_{j}p_{j}\mu^{j}_{B_{1}A^{\prime}_{1}}\otimes\tilde{\mu}^{j}_{B_{2}A^{\prime}_{2}}$, where $p_{j}\geq 0$ and $\sum_{j}p_{j}=1$. Let $\rho_{A}=\sum_{k}q_{k}\rho^{k}_{A_{1}}\otimes\tilde{\rho}^{k}_{A_{2}}$, where $q_{k}\geq 0$ and $\sum_{k}q_{k}=1$, be another separable state. Then
@@
$\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ $=\operatorname{Tr}_{A}[\mu_{BA}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T})]$ (6)
$=\sum_{j,k}p_{j}q_{k}\sigma^{j,k}_{B_{1}}\otimes\tilde{\sigma}^{j,k}_{B_{2}}.$
where $\sigma^{j,k}_{B_{1}}\coloneqq\operatorname{Tr}_{A_{1}}\{\mu^{j}_{B_{1}A_{1}}[\operatorname{\mathds{1}}_{B_{1}}\otimes(\rho^{k}_{A_{1}})^{T}]\}$, and $\tilde{\sigma}^{j,k}_{B_{2}}\coloneqq\operatorname{Tr}_{A_{2}}\{\tilde{\mu}^{j}_{B_{2}A_{2}}[\operatorname{\mathds{1}}_{B_{2}}\otimes(\tilde{\rho}^{k}_{A_{2}})^{T}]\}$. This proves that $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ is a separable state whenever $\mu_{BA^{\prime}}$ and $\rho_{A}$ are separable, therefore condition 2 of Theorem 3 is satisfied. This implies that it is possible to construct a CDRT from the set of separable states which coincides with SEP [37, 38] as noted above. Once again, since the identity matrix is separable, the maximally mixed state is free, as stated in Corollary 4.
§ VIII.  Example: resource theory of NPT entanglement
Looking again at the resource theory of entanglement, one notices that all separable states have positive partial transpose [39, 40]. One can thus consider as free states the bipartite states with positive partial transpose. This set is strictly larger than the set of separable states due to bound entanglement [41], which is of great relevance in quantum information [39, 40, 41, 42, 43, 44, 45]. Since the set of PPT states contains separable states, it automatically satisfies condition 1 of Theorem 3. Let $\mu_{BA^{\prime}}$ and $\rho_{A}$ be free states. We write $(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}$ as
$(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}$ $=\operatorname{Tr}_{A}(\mu_{B\mathord{:}A}^{T_{B_{2}}}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T}))$ (7)
$=\operatorname{Tr}_{A}[\mu_{BA}^{T_{B_{2}A_{2}}}(\operatorname{\mathds{1}}_{B}\otimes\rho_{A}^{T_{A_{1}}})].$
Since $\rho_{A}$ has positive partial transpose, there exists and eigenbasis {|$\Set{\Ket{\varphi^{j}_{A}}}$ and positive eigenvalues $\lambda_{j}$ such that $\rho_{A}^{T_{A_{1}}}=\sum_{j}\lambda_{j}|\varphi^{j}_{A}\rangle\!\langle\varphi^{j}_{A}|$. This implies that, for all |$\Ket{\psi_{B}}$,
$\Braket{\psi_{B}}{(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}}{\psi_{B}}=$ (8)
$\sum_{j}\lambda_{j}\Braket{\psi_{B}\otimes\varphi_{A}^{j}}{\mu_{BA}^{T_{B_{2}A_{2}}}}{\psi_{B}\otimes\varphi_{A}^{j}}\geq 0,$
where the last inequality follows from the fact that $\mu_{BA}^{T_{B_{2}A_{2}}}\geq 0$, as it is free, and $\lambda_{i}>0$. This is proves that $d_{A}(\mu_{B\mathord{:}A^{\prime}}*\rho_{A})^{T_{B_{2}}}\geq 0$, and that the set of states with positive partial transpose is closed under link product. In this case too, we have shown that the set of free states satisfies the conditions of Theorem 3. The CDRT constructed coincides with the theory of NPT [42, 43, 44, 45].
§ IX.  Conclusions

### v2

@@ Abstract
Many resource theories share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state. In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories. We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations. Moreover, we define resource measures, a complete family of monotones, and conversion distances in such resource theories.
@@
The recent booming interest in quantum computing, communication, and information originates from the understanding that quantumness allows us to outperform the classical world, which means that quantum objects are resources. The new technological developments, which are now becoming commercially available, demand a rigorous and thorough understanding of quantum resources, in order to quantify them precisely, and optimize their use. This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Resource theories are based on the simple observation that agents in a laboratory can only perform a subset of all possible deterministic quantum operations due to the various constraints on the agents and the systems involved. For instance, when two agents, Alice and Bob, are physically separated, they may not have a reliable means to exchange quantum systems, limiting them to local operations and classical communication (LOCC) [5, 6, 7]. Such allowed operations are called free.
@@
While free states are typically defined as those states that can be prepared with free operations, there are situations where it is beneficial to construct free operations starting from a set of free states. As an example, Alice and Bob may quickly realize that they can create all and only separable states and wonder what free operations are compatible with their set of free states. There are different ways to construct free operations starting from free states [1]. However, the minimal condition is that one cannot generate a resource by applying a free operation on a free state in a complete sense, i.e., even if applied only to part of a bipartite state. Operations defined with such a construction are known as completely resource non-generating (CRNG) operations [1]. CRNG operations are of particular interest because they are the largest set of free operations compatible with a given set of free states, and they can be used to derive lower or upper bounds in any resource theory with the same set of free states. For example, separable and positive partial transpose operations have been used to overcome the challenging characterization of LOCC operations [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. These extensions of the resource theory of entanglement are known as the resource theories of separable entanglement (SEP) and non positive partial transpose entanglement (NPT), respectively. In other settings, such as in the resource theory of magic [28, 29, 30, 31, 32], imaginarity [33, 34, 35], non-negativity of quantum amplitudes [47], the CRNG operations are even more relevant because they coincide with the free operations.
Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state. We call resource theories that exhibit this property Choi-defined resource theories. This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization.
Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states. In this letter, we answer this question by providing necessary and sufficient conditions on the set of free states. Moreover, we show that when such a construction is possible, the resulting set of free operations coincides with the set of CRNG operations associated with those free states. These results provide a constructive definition of CRNG operations and generalize what has been already observed in the resource theories of magic, imaginarity, NPT, SEP, and non-negativity of quantum amplitudes. Next, we introduce resource measures, a complete family of monotones, and conversion distances in CDRTs, which can all be computed with conic linear programs (CLP) [48, 49] whenever the set of free states is convex and closed, or even with semidefinite programs (SDP) [50]. Lastly, we present the resource theory of asymmetry [51], for which the correspondence between the free states and the Choi matrices of the free channels was not already known, and we also characterize which resource theories of asymmetries are CDRTs.
@@
One of the constraints of the resource theory of bipartite entanglement is that the two physically separated agents cannot exchange quantum systems. This may appear to conflict with the condition that the swap channel is a free operation. However, in this scenario, the systems in question are bipartite systems [2], i.e., $S_{1}\coloneqq A_{1}B_{1}$ and $S_{2}\coloneqq A_{2}B_{2}$. The physical separation is, as usual, between systems labeled with $A$ and systems labeled with $B$. The swap channel simply swaps the order of the two bipartite systems $S_{1}$ and $S_{2}$, i.e., $A_{1}B_{1}A_{2}B_{2}\leftrightarrow A_{2}B_{2}A_{1}B_{1}$. If we organize the systems according to the physical separation, the action of the swap channel is $(A_{1}A_{2})(B_{1}B_{2})\to(A_{2}A_{1})(B_{2}B_{1})$. Therefore, the two agents are locally exchanging the order of their own quantum systems and there is no exchange of quantum systems between them.
@@
Note that here one needs to know which system is the input of the original channel and which is the output. Clearly, if such information is missing, a matrix $M_{BA}$ could be associated with different linear maps, see the Supplemental Material (SM) for examples.
Observe that, if we divide the Choi matrix $M_{B\mathord{:}A^{\prime}}$ of the channel $\mathcal{M}_{A\to B}$ by the dimension $d_{A}$ of $A$, we obtain a quantum state $\mu_{B\mathord{:}A^{\prime}}$ such that
@@
It is a natural question to characterize when, given a set of free states, the construction of a CDRT is allowed. In every well-defined resource theory, the identity is a free operation. As a consequence, its renormalized Choi state $\frac{1}{d_{A}}\Phi_{AA^{\prime}}$ is free in every CDRT.
Remark 2.
The renormalized Choi state is not the maximally entangled state in the resource theories of entanglement, SEP entanglement, or NPT entanglement. Indeed, in these resource theories, each system is a pair of pysical systems [2], i.e., $S\coloneqq AB$, and $S^{\prime}\coloneqq A^{\prime}B^{\prime}$ and the spatial separation is between systems $AA^{\prime}$ and $BB^{\prime}$. Let {|$\Set{\Ket{a}_{A}}$ and {|$\Set{\Ket{b}_{B}}$ be orthogonal basis for $A$ and $B$, respectively, then
$\frac{1}{d_{S}}\Phi_{SS^{\prime}}=\frac{1}{d_{A}d_{B}}\sum_{a,b}|ab\rangle\!\langle ab|_{AB}\otimes|ab\rangle\!\langle ab|_{A^{\prime}B^{\prime}}.$ (4)
If the systems are reorganized according to the pysical separation, the renormalized Choi state becomes
$\frac{1}{d_{A}d_{B}}\sum_{a}|aa\rangle\!\langle aa|_{AA^{\prime}}\otimes\sum_{b}|bb\rangle\!\langle bb|_{BB^{\prime}}.$ (5)
This shows that the state $\frac{1}{d_{S}}\Phi_{SS^{\prime}}$ is separable with respect to the spatial separation between $AA^{\prime}$ and $BB^{\prime}$. Therefore, it is free in the resource theories mentioned above, so it is not the maximally entangled state on $(AA^{\prime})(BB^{\prime})$.
Requesting that resource theories be closed under the sequential composition of free channels (Figure 1), we find a less trivial condition for the set free states. In Ref. [52, 53], the authors introduced an operation between bipartite matrices, called link product, defined as
@@
where $\cdot^{T_{B}}$ is the partial transpose on system $B$. This operation translates sequential composition of channels into an operation between Choi matrices. Thanks to the link product, we can restate the golden rule of resource theories as a condition on the set of free states. That is, the condition that $\mathcal{M}_{A\to B}(\rho_{A})$ be free whenever $\mathcal{M}_{A\to B}$ and $\rho_{A}$ are free becomes that $d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$ be a free state if $\rho_{A}$ and $\mu_{B\mathord{:}A^{\prime}}$ are free states and $\mu_{B\mathord{:}A^{\prime}}$ is the renormalized Choi matrix of a quantum channel.
We now state our first main result.
@@
Notably, this theorem provides an easy construction for CRNG operations when the set of free states satisfies the conditions of Theorem 3. For example, the set of separable states satisfies the conditions of Theorem 3, as we show in the SM. Therefore, Theorems 3 and 5 imply that the CRNG operations are those that have a separable Choi matrix, i.e., the separable operations. Note that in general CRNG operations are not CD operations. Indeed, the CRNG operations in the resource theory of athermality are the Gibbs-preserving operations (see, e.g., Ref. [56]), but their renormalized Choi matrices are not free states.
§ V.  Quantifying resources in CDRTs
CDRTs are constructed from a given set of free states that satisfies the conditions of Theorem 3. As a consequence, they inherit all the resource measures that depend only on the set of free states. Examples of such resource measures are the robustness, the generalized robustness, and all the ones based on quantum divergences (see Ref. [1] for a review). Among these, the max-relative entropy [57] is of particular interest in the context of CDRTs. It is defined as
$D_{\max}(\rho_{A}\|\sigma_{A})=\log\min\Set{\lambda\mid\rho_{A}\leq\lambda\sigma_{A}},$ (7)
when $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\sigma_{A}$ and +$\infty$+ ∞ otherwise. The extension of the max-relative entropy to quantum channels [58, 59] satisfies [60]
$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=D_{\max}(\mu_{BA}\|\nu_{BA}),$ (8)
where, $\mu_{BA}$ and $\nu_{BA}$ are the renormalized Choi matrices of $\mathcal{M}_{A\to B}$ and $\mathcal{N}_{A\to B}$, respectively. In a CDRT, this property allows us to define resource measures for states [1] and channels [61] that depend uniquely on the set of free states:
$D_{\max}^{\mathfrak{F}}(\rho_{A})=\inf_{\omega_{A}\in\mathfrak{F}(A)}D_{\max}(\rho_{A}\|\omega_{A}),$ (9)
$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\begin{subarray}{c}\omega_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\omega_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D_{\max}(\mu_{BA}\|\omega_{BA}),$
where $\mathfrak{F}(A)$ denotes the set of free states of system $A$. Both these resource measures are finite in every CDRTs because for every $\rho_{A}$, $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\operatorname{\mathds{1}}_{A}$ and $\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A}$ is free. The same is not true for all resource theories. Moreover, these resource measures can be computed with a CLP whenever the set of free states is convex and closed, and with a SDP whenever the membership problem for the set of free states is an SDP (see the SM).
As a direct consequence of the results in Ref. [60], the resource measures in Eq. (9) satisfy
$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B}(\rho_{A}))\leq D_{\max}^{\mathfrak{F}}(\rho_{A})+D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})$ (10)
for all states $\rho_{A}$ and channels $\mathcal{M}_{A\to B}$. For the reasons mentioned above, this inequality, which is valid for all resource theories, is particularly meaningful in the context of CDRTs. When we interpret $D^{\mathfrak{F}}$ as the value of a resource, it implies that value of the output of a quantum operation applied to a quantum state is always less than the sum of the values of the operation and the input state. Equivalently, for every $\rho_{A}$, $\sigma_{B}$, and $\mathcal{M}_{A\to B}$ such that $\sigma_{B}=\mathcal{M}_{A\to B}(\rho)$, the cost of preparing $\sigma_{B}$ is always less or equal than the cost of preparing $\rho_{A}$ and then converting it to $\sigma_{B}$ with the channel $\mathcal{M}_{A\to B}$. An inequality similar to the one in Eq. (10) was derived in Ref. [62]. The divergences that appear in that inequality are the relative entropy and the regularized channel relative entropy. Instead, in our case, only the max-relative entropy is used, and no regularization is needed.
§ VI.  Resource conversion in CDRTs
Where the formalism of CDRTs provides the greatest insight is in all quantities that are connected with resource conversion or manipulation via free operations. Indeed at the core of every resource theory there is a preorder of resources defined as $\rho_{A}\succ\sigma_{B}$ if there exists a free channel $\mathcal{M}_{A\to B}$ such that $\sigma_{B}=\mathcal{M}_{A\to B}(\rho_{A})$. In a CDRT, this condition is equivalent to requiring the existence of a free states $\mu_{BA}$, such that $\operatorname{Tr}_{B}\mu_{BA}=\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A}$ and $\sigma_{B}=d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}$.
As a consequence, many quantities or results that involve free channels can be redefined in terms of the link product of free states. A first example is the complete family of monotones defined in Ref. [61] for convex and closed resource theories, which in the case of a convex and closed CDRT can be expressed as
$f_{\tau_{B}}(\rho_{A})=\max_{\begin{subarray}{c}\omega_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\omega_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}\operatorname{Tr}_{BA}[\omega_{BA}(\tau_{B}\otimes\rho_{A}^{T})].$ (11)
Being a complete family of monotones, it satisfies $f_{\tau_{B}}(\rho_{A})\geq f_{\tau_{B}}(\sigma_{B})$ for all density matrices $\tau_{B}$ if and only if $\rho_{A}$ can be converted to $\sigma_{B}$ with a free operation. The advantage of this approach is that the optimization is over states and not channels. In particular, the quantity above can be computed with a CLP (see the SM).
As a second example, we present conversion distances in CDRTs. These distances estimate how close to a target state an initial state can be converted by using only free operations. In a CDRT, they satisfy
$D(\rho_{A}\to\sigma_{B})$ $=\inf_{\mathcal{M}_{A\to B}\in\mathfrak{F}(A\to B)}D(\sigma_{B},\mathcal{M}_{A\to B}(\rho_{A}))$ (12)
@@
$=\inf_{\begin{subarray}{c}\mu_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\mu_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D(\sigma_{B},d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}),$
where $D(\cdot,\cdot)$ is a function on quantum states that satisfies the data processing inequality [1]. Once again, the minimization is only over states and this can simplify the computation of the conversion distances. In particular, many metrics, such as the trace distance or the purified distance, can be computed with SDPs [50]. Consequently, the conversion distances based on such metrics are SDPs too whenever the membership problem for the set of free states can be expressed as an SDP, e.g., in the case of the resource theories of imaginarity and PPT.
§ VII.  Examples of CDRTs
As mentioned in the introduction, known results about the resource theories of SEP, PPT, magic, imaginarity, and non-negative quantum amplitudes imply that these resource theories are CDRTs and we show in the SM that their sets of free states satisfy the conditions of Theorem 3.
Of more interest are those resource theories where the identification of free operations with CD operation has not been already established. For example, to the best of our knowledge, there is no result yet in this direction for the resource theory of asymmetry. In this section, we derive under what conditions resource theories of asymmetry [51, 14] are CDRTs.
In the resource theory of asymmetry, two parties, Alice and Adam have access to the same quantum systems but they do not share a reference frame. The relation between reference frames is described by an element $g$ of a compact group $G$ identified by the symmetries of the problem. If Alice and Adam do not know how their reference frames are related, then Alice can prepare only states for Adam that are $G$-invariant, that is:
$\rho_{A}\,\text{is free}\,\Leftrightarrow\mathcal{U}_{g}^{A}(\rho_{A})=\rho_{A},\,\forall g\in G,$ (13)
where $\mathcal{U}_{g}^{A}(\rho_{A})\coloneq U^{A}_{g}\rho_{A}(U^{A}_{g})^{\dagger}$, and $U_{g}$ is the unitary representation of $g$ on $A$. Similarly, if Adam asks Alice to perform a deterministic quantum operation, then she can only perform $G$-covariant channels:
$\mathcal{M}_{A\to B}\,\text{is free}\,\Leftrightarrow\mathcal{U}^{B}_{g}\circ\mathcal{M}_{A\to B}\circ(\mathcal{U}_{g}^{A})^{\dagger}=\mathcal{M}_{A\to B}$ (14)
for all $g\in G$.
Every different finite or compact Lie group $G$ generates a different resource theory, and the following proposition characterizes the ones that are CDRTs.
Proposition 6.
A resource theory of asymmetry is a CDRT if and only if the renormalized Choi state is free.
We point out that this is equivalent to asking that all the unitary representations are real (see the SM for details). An example of such resource theory is the parity resource theory generated by the group $\mathbb{Z}_{2}$ [51].
§ VIII.  Conclusions